I'm thoroughly stuck on this trig problem: [(cosine of theta times cotangent of theta) over (1 minus sin of theta)] minus 1 equals cosecant of theta. I've tried converting everything to sines and cosines, and that didn't work, and I tried everything I could think of with the fundamental identities and nothing worked there either. Any help would be greatly appreciated.
Found 2 solutions by Alan3354, richard1234:
Answer by Alan3354(69443)

cos*cot/(1 - sin) - 1 = csc
---------
You didn't say what you want to do.
-----
(cos^2/sin)/(1 - sin) - 1 = csc
((1-sin^2)/sin)/(1-sin) - 1 = csc
(1 + sin)/sin - 1 = csc
1/sin + sin/sin - 1 = csc
1/sin = csc
csc = csc

Answer by richard1234(7193)